# Emergence of a hidden-order phase well below the charge density wave transition in a topological Weyl semimetal (TaSe4)2I

### v1

@@ D First-principles calculations
To further uncover the origin of the hidden phase transition, we performed first-principles band structure calculations for the three structures of TSI2 to analyze the change near the Fermi level. The plots are summarized in Fig. 5. The band structures for the room-temperature phase $I422$ and mid-temperature (i.e. $T^{*}<T<T_{CDW}$) structures $I4$ look identical, as shown in Fig. 5 (a). As one can see, DFT calculations predict a semi-metallic nature, and we could not find any noticeable change in the band structure for these two structures. At the Fermi level, the contribution comes mostly from Ta-$d$ orbitals (not shown). TSI2 consists of indefinite TaSe4 chains with each Ta ion being sandwiched between two consecutive $Se_{4}^{4-}$ rectangles. The equivalent number of $d$ electrons is $\frac{1}{2}$ in each Ta atom, and the corresponding band filling of Ta-$d$ orbitals here is $\frac{1}{4}$[gressier1984electronic]. This partial filling of bands facilitates Peierl’s distortion.
@@
Nevertheless, the CDW phase in (TaSe4)2I emerges below $T_{CDW}$ ($\sim$263K), the subsequent noticeable tetramerization involving Ta atoms along the chain-axis could be captured much below $T_{CDW}$[van2001structure].
@@
We further analyze the temperature-dependent phase transition in TSI2 by investigating the point-group symmetry of the three crystal structures using the CELLSUB module [bb2] of Bilbao crystallographic server [bb1]. There can be three different pathways to $I422$ $\rightarrow$ $C2$ transition via an intermediate phase, which can be $F222$, $I222$, or $I4$ as shown in Fig. 6. However, as in our experiments we obtained the $I4$ structure in the intermediate temperature range, this eliminates the possibility first two intermediate phases. During the transition, the point group symmetry lowers from $D_{4}$ in the $I422$ structure to $C_{2}$ in $C_{2}$ phase via $C_{4}$ of $I4$ structure. Further analysis using the AMPLIMODES module [amp1, amp2] of the Bilbao Crystallographic Server reveals that in the case of $I422\rightarrow I4$ transition, it is $\sim$0.007 Å atomic displacements of 8cc Wycoff positions of Se atoms which causes the lowering of symmetry from $D_{4}$ to $C_{4}$. This strongly suggests a lattice-driven origin for the observed hidden-order transition near $T^{*}$ and reinforces the idea that subtle structural instabilities, coupled to electronic degrees of freedom, underpin the emergence of a gapped, correlated ground state well below TCDW.
Figure 6: (a) Evolution of crystal structure analyzed using the CELLSUB module [bb2] of Bilbao crystallographic server [bb1]. The green arrow indicates the energetically feasible pathway among the three possible ways for TSI2.(b-c) Schematic representation of TaSe4 chains with different tetramerized structures of periodic LLSS(pink octahedra) and SSLL(blue octahedra) units, respectively. The (1,1,0) plane is marked by a dashed line. Two consecutive (1,1,0) planes sandwich TaSe4 chains with an alternate sequence of LLSS and SSLL bonding sequence.
@@ III Discussion
The study of charge density wave (CDW) dynamics and collective excitations in (TaSe4)2I has garnered considerable attention due to its complex band topology[Shi2021, Zahid_Nat.Phy, li2021typeIIIWeyl] and recent claims of axion-like excitations [AxionGooth2019]. Notably, the realization of a massive phason mode via a Higgs-like mechanism has been proposed to occur exclusively below the secondary transition temperature $T^{*}$ [bae2025trSTM, kim2023observation]. A recent synchrotron study on TSI2 argued the possibility of Ta-tetramerization and monoclinic distortion only below 120 K[van2001structure]. This was later anticipated in another recent X-ray experiment [Structure_PhysRevLett.87.015502], further corroborating a symmetry-lowered ground state distinct from the primary CDW phase.
In this context, low-frequency resistance noise provides a powerful window for tracing the underlying fluctuations and emergent order. In broken-symmetry systems exhibiting long-range order, low-frequency fluctuations are often governed by the relaxation dynamics of carriers and defects with finite lifetimes [kogan1984low, RevModPhysWeissman, CuerrentOpinion_Raychaudhuri]. In non-magnetic conductors, $1/f$ noise typically arises from slow fluctuations in either the carrier density or mobility [gruner2018density]. Moreover, since the resistance $R$ of a non-magnetic conductor depends on both carrier mobility $\mu$ and carrier density $n$, the normalized variance in resistance fluctuations can be expressed as [Hooge1994_1byf_noise, hooge1981experimental, RevModPhysWeissman, Kalimuddin.PRL]:
@@
In our recent study [Kalimuddin.PRL], we argued that fluctuations in the condensed carrier density ($n_{c}$) are transferred to normal carriers ($n_{n}$) using a two-fluid analogy. At high temperatures, resistance fluctuations are primarily governed by amplitude modes near $T_{\mathrm{CDW}}$, whereas at lower temperatures near $T^{*}$, the low density of thermally activated quasiparticles points to an increasing contribution from phase fluctuations. The emergence of strong correlations in noise statistics near $T^{*}$—evident through non-Gaussian probability distributions and enhanced second spectra—further indicates the possible presence of a secondary, symmetry-breaking transition distinct from the conventional CDW order.
It is also noteworthy that our temperature-dependent single-crystal Laue diffraction reveals subtle but distinct anomalies in the lattice parameters (a and c), unit-cell volume, Se–Se dimer bond length, and the dihedral angle between consecutive $Se_{4}$} planes near the $T^{*}$ (see Fig. S1 of the SM[supply]). These variations indicate a symmetry-lowering distortion concurrent with the onset of enhanced low-frequency noise and thermopower anomalies, suggesting a lattice-driven hidden-order phase. This observation is consistent with previous low-temperature X-ray diffraction studies in (TaSe4)2I[van2001structure, Structure_PhysRevLett.87.015502].
@@ V Acknowledgements
This work was supported by (i) CSIR. Grant Number: 03/ 1511/23/EMR-II. (ii) Department of Science and Technology, Government of India (CRG/2023/001100). The authors thank Riju Pal and Prof. Atindra Nath Pal from SNBNCBS for the technical help. S. Chatterjee thanks Dr. Nitish Mathur for the helpful discussions. Sk Kalimuddin thanks IACS for PhD fellowship. TD and SD acknowledge UGC, India, for support from fellowship. SKP thanks NFSG grant from BITS-Pilani, Dubai campus which supported this research.
@@ Abstract
The Seebeck coefficient (S), a fundamental parameter in thermoelectric materials, quantifies the voltage generated in response to a temperature gradient across a material. Accurate measurement of this coefficient as a function of temperature is crucial for evaluating and optimizing materials for thermoelectric applications. Here, we employed the steady-state method similar to our previous work[Sruthi2022], where we ensured a steady temperature gradient across the sample and measured the resulting thermoelectric voltage using a nano-voltmeter(Keithley 2182A). The Seebeck coefficient is determined from the ratio of the voltage difference to the temperature difference[Patel04072017]. The experimental setup for the Seebeck coefficient has been shown in Figure S3(a).
@@ A Calibration of Seebeck setup
The absolute calibration of the Seebeck has been done using a high-quality (99.99 % pure ) Nickel piece and a platinum wire of purity 99.99 %. Our experimentally measured Seebeck coefficient over a temperature window of 80-375 K matches well with reference data for both Nickel[Seebeck_Nickel] and Platinum[Seebeck_Platinum] respectively. The experimental Seebeck data vs reference data plot for Ni and Pt wire are shown in Figure S3(b-c) respectively.
Figure S3: Thermoelectric power setup and calibration: (a) Liquid nitrogen-based cryogenic insert for thermo-emf measurement. (b) (top) Image of the sample holder with (TaSe4)2I mounted on it, and (bottom) Wiring configuration for local heating and measurement of thermo-electric voltage. (b) Experimental Seebeck(purple) vs reference data[Seebeck_Nickel] (blue) for Nickel. (c) Experimental Seebeck(purple) vs reference data[Seebeck_Platinum] (blue) for platinum wire.

### v2

@@ D First-principles calculations
To further uncover the origin of the hidden phase transition, we performed first-principles band structure calculations for the three structures of TSI2 to analyze the change near the Fermi level. The plots are summarized in Fig. 5. The band structures for the room-temperature phase $I422$ and mid-temperature (i.e. $T^{*}<T<T_{CDW}$) structures $I4$ look identical, as shown in Fig. 5 (a). As one can see, DFT calculations predict a semi-metallic nature, and we could not find any noticeable change in the band structure for these two structures. At the Fermi level, the contribution comes mostly from Ta-$d$ orbitals (not shown). TSI2 consists of indefinite TaSe4 chains with each Ta ion being sandwiched between two consecutive $Se_{4}^{4-}$ rectangles. The equivalent number of $d$ electrons is $\frac{1}{2}$ in each Ta atom, and the corresponding band filling of Ta-$d$ orbitals here is $\frac{1}{4}$[25]. This partial filling of bands facilitates Peierl’s distortion.
@@
Nevertheless, the CDW phase in (TaSe4)2I emerges below $T_{CDW}$ ($\sim$263K), the subsequent noticeable tetramerization involving Ta atoms along the chain-axis could be captured much below $T_{CDW}$[65].
@@
We further analyze the temperature-dependent phase transition in TSI2 by investigating the point-group symmetry of the three crystal structures using the CELLSUB module [4] of Bilbao crystallographic server [3]. There can be three different pathways to $I422$ $\rightarrow$ $C2$ transition via an intermediate phase, which can be $F222$, $I222$, or $I4$ as shown in Fig. 6. However, as in our experiments we obtained the $I4$ structure in the intermediate temperature range, this eliminates the possibility first two intermediate phases. During the transition, the point group symmetry lowers from $D_{4}$ in the $I422$ structure to $C_{2}$ in $C_{2}$ phase via $C_{4}$ of $I4$ structure. Further analysis using the AMPLIMODES module [48, 51] of the Bilbao Crystallographic Server reveals that in the case of $I422\rightarrow I4$ transition, it is $\sim$0.007 Å atomic displacements of 8cc Wycoff positions of Se atoms which causes the lowering of symmetry from $D_{4}$ to $C_{4}$. This strongly suggests a lattice-driven origin for the observed hidden-order transition near $T^{*}$ and reinforces the idea that subtle structural instabilities, coupled to electronic degrees of freedom, underpin the emergence of a gapped, correlated ground state well below TCDW.
Figure 6: (a) Evolution of crystal structure analyzed using the CELLSUB module [4] of Bilbao crystallographic server [3]. The green arrow indicates the energetically feasible pathway among the three possible ways for TSI2.(b-c) Schematic representation of TaSe4 chains with different tetramerized structures of periodic LLSS(pink octahedra) and SSLL(blue octahedra) units, respectively. The (1,1,0) plane is marked by a dashed line. Two consecutive (1,1,0) planes sandwich TaSe4 chains with an alternate sequence of LLSS and SSLL bonding sequence.
@@ III Discussion
The study of charge density wave (CDW) dynamics and collective excitations in (TaSe4)2I has garnered considerable attention due to its complex band topology[58, 41, 39] and recent claims of axion-like excitations [23]. Notably, the realization of a massive phason mode via a Higgs-like mechanism has been proposed to occur exclusively below the secondary transition temperature $T^{*}$ [6, 32]. A recent synchrotron study on TSI2 argued the possibility of Ta-tetramerization and monoclinic distortion only below 120 K[65]. This was later anticipated in another recent X-ray experiment [18], further corroborating a symmetry-lowered ground state distinct from the primary CDW phase.
In this context, low-frequency resistance noise provides a powerful window for tracing the underlying fluctuations and emergent order. In broken-symmetry systems exhibiting long-range order, low-frequency fluctuations are often governed by the relaxation dynamics of carriers and defects with finite lifetimes [33, 69, 52]. In non-magnetic conductors, $1/f$ noise typically arises from slow fluctuations in either the carrier density or mobility [26]. Moreover, since the resistance $R$ of a non-magnetic conductor depends on both carrier mobility $\mu$ and carrier density $n$, the normalized variance in resistance fluctuations can be expressed as [28, 29, 69, 30]:
@@
In our recent study [30], we argued that fluctuations in the condensed carrier density ($n_{c}$) are transferred to normal carriers ($n_{n}$) using a two-fluid analogy. At high temperatures, resistance fluctuations are primarily governed by amplitude modes near $T_{\mathrm{CDW}}$, whereas at lower temperatures near $T^{*}$, the low density of thermally activated quasiparticles points to an increasing contribution from phase fluctuations. The emergence of strong correlations in noise statistics near $T^{*}$—evident through non-Gaussian probability distributions and enhanced second spectra—further indicates the possible presence of a secondary, symmetry-breaking transition distinct from the conventional CDW order.
It is also noteworthy that our temperature-dependent single-crystal Laue diffraction reveals subtle but distinct anomalies in the lattice parameters (a and c), unit-cell volume, Se–Se dimer bond length, and the dihedral angle between consecutive $Se_{4}$} planes near the $T^{*}$ (see Fig. S1 of the SM[56]). These variations indicate a symmetry-lowering distortion concurrent with the onset of enhanced low-frequency noise and thermopower anomalies, suggesting a lattice-driven hidden-order phase. This observation is consistent with previous low-temperature X-ray diffraction studies in (TaSe4)2I[65, 18].
@@ V Acknowledgements
This work was supported by CSIR-Human Resource Development Group (HRDG) Grant No. (03/ 1511/23/EMR-II) and Department of Science and Technology, Government of India (CRG/2023/001100). The authors thank Riju Pal and Prof. Atindra Nath Pal from SNBNCBS for the technical help. S. Chatterjee thanks Dr. Nitish Mathur for the helpful discussions. Sk Kalimuddin thanks IACS for PhD fellowship. TD and SD acknowledge UGC, India, for support from fellowship. SKP thanks NFSG grant from BITS-Pilani, Dubai campus which supported this research.
@@ Abstract
The Seebeck coefficient (S), a fundamental parameter in thermoelectric materials, quantifies the voltage generated in response to a temperature gradient across a material. Accurate measurement of this coefficient as a function of temperature is crucial for evaluating and optimizing materials for thermoelectric applications. Here, we employed the steady-state method similar to our previous work[54], where we ensured a steady temperature gradient across the sample and measured the resulting thermoelectric voltage using a nano-voltmeter(Keithley 2182A). The Seebeck coefficient is determined from the ratio of the voltage difference to the temperature difference[49]. The experimental setup for the Seebeck coefficient has been shown in Figure S3(a).
@@ A Calibration of Seebeck setup
The absolute calibration of the Seebeck has been done using a high-quality (99.99 % pure ) Nickel piece and a platinum wire of purity 99.99 %. Our experimentally measured Seebeck coefficient over a temperature window of 80-375 K matches well with reference data for both Nickel[11] and Platinum[46] respectively. The experimental Seebeck data vs reference data plot for Ni and Pt wire are shown in Figure S3(b-c) respectively.
Figure S3: Thermoelectric power setup and calibration: (a) Liquid nitrogen-based cryogenic insert for thermo-emf measurement. (b) (top) Image of the sample holder with (TaSe4)2I mounted on it, and (bottom) Wiring configuration for local heating and measurement of thermo-electric voltage. (b) Experimental Seebeck(purple) vs reference data[11] (blue) for Nickel. (c) Experimental Seebeck(purple) vs reference data[46] (blue) for platinum wire.